OpenStudy (anonymous):

the formula f= 9/5c +32 gives the fahrenheit f temperature in terms of celsius c temperature when the farenheit temperture is 50

OpenStudy (anonymous):

A.90 B.82 C.18 D.10 EXPLAIN

OpenStudy (anonymous):

when fahrenheit temp is 50, just plug in 50 in the place of F in your formula. Therefore, 50 = (9/5) C + 32 (9/5) C = 50 - 32 = 18 C = 18*5/9 = 10 Therefore, temperature in 10 degree celcius

OpenStudy (anonymous):

OK SO THE ONLY THING I HAD TO DO WAS PLUG IN THE NUMBER 50

OpenStudy (anonymous):

AND THANKS